# PerfBench: Can Agents Resolve Real-World Performance Bugs?

### v2

@@ Abstract
Performance bugs are inefficiencies in software that waste computational resources without causing functional failures, making them particularly challenging to detect and fix. While recent advances in Software Engineering agents have shown promise in automated bug fixing, existing benchmarks primarily focus on functional correctness and fail to evaluate agents’ abilities to identify and resolve non-functional issues like performance bugs. We introduce PerfBench, a benchmark comprising 81 real-world performance bug-fixing tasks from popular .NET repositories on GitHub. Unlike existing benchmarks that rely on pre-existing test suites, PerfBench features a novel evaluation harness that allows agents to generate their own performance benchmarks and validates fixes by comparing execution metrics collected for developer fix and agent fix. Each task in PerfBench is derived from actual developer fixes linked to performance-related issues, which are then verified by human experts, ensuring real-world relevance. Our evaluation reveals that current state-of-the-art coding agents struggle with performance optimization tasks, with baseline OpenHands agent achieving only a $\sim$3% success rate on our benchmark. We develop OpenHands-Perf-Agent, which incorporates performance-aware tooling and instructions and achieves a $\sim$20% success rate on the benchmark. We show that by ensuring the agent has proper instructions to benchmark its changes and tooling for benchmark output processing, we can improve the agent performance significantly, but room for improvement still remains. PerfBench provides a challenging test set for furthering the capabilities of agents in fixing performance issues.
I Introduction
@@
Performance bugs represent a unique class of software defects that impact the application’s efficiency without causing any functional failures. Unlike traditional bugs that manifest as crashes or incorrect outputs, performance bugs silently waste computational resources, increase latency as well as costs [1, 2]. These inefficiencies are particularly problematic in cloud applications where resource consumption directly translates to computational costs, and can also impact end-user experience due to problems such as increased latency.
@@
The rise of Software Engineering agents has revolutionized automated software engineering, with proprietary systems like Claude Code [3], Copilot Agent [4], Windsurf [5], Devin [6], etc. as well as open-source agents such as SWE-agent [7], and OpenHands [8] demonstrating impressive capabilities in fixing functional bugs as well as other software engineering tasks such as test generation, code search, feature development, etc. However, current bug fixing benchmarks for evaluating these agents, such as SWE-bench [9], focus exclusively on functional correctness. This leaves a significant gap in understanding how well these agents can handle non-functional bugs such as performance or security bugs.
@@
II Background and Related Work
@@
We discuss work closely related to performance bug detection and repair, benchmarks for code generation, and LM agents for software engineering.
II-A Performance Bugs in Software Systems
@@
Performance bugs represent a unique class of software defects that impact efficiency without causing functional failures. They tend to be harder to detect  [1, 2, 10, 11] and fix  [12, 13] than functional bugs. Due to being hard to detect and not causing outright failures, these bugs can often go undetected for long periods of time  [10, 11]. As a result, better tool support is needed to fix performance bugs. While performance profiling have been developed to identify these issues [10, 1, 14], these bugs require significant manual analysis to translate findings into fixes.
II-B Automated Program Repair for Performance Issues
@@
Traditional automated program repair (APR) techniques focus primarily on functional correctness using test suites or logic assertions as specifications. GenProg [15] uses genetic programming to generate fixes that pass test cases, while pattern-based approaches like PAR [16] leverage manually created fix templates. Similarly, CapGen [17] generates patches at finer granularity using context-aware prioritization, and VarFix [18] extends GenProg through variational execution [19]. However, these traditional APR approaches face significant challenges when applied to performance bugs. Performance issues require different validation mechanisms than functional correctness, as they must demonstrate measurable improvements in efficiency metrics rather than passing test cases. Additionally, performance bugs often involve subtle inefficiencies that require domain-specific knowledge about algorithms, data structures, and system behavior that is difficult to encode in traditional search-based approaches. Not only that, performance bugs manifest in various forms including algorithmic inefficiencies, memory leaks, excessive allocations, and suboptimal API usage patterns. As a result, for a fix approach to be useful, it would need to be general enough to fix a wide-range of issues. However, many existing approaches target specific kinds of issues such as repeated computations [20], software misconfigurations [21], loop inefficiencies [12]. Recent work in model training has explored performance-specific repair. Studies like DeepDev-PERF [22], RAPGen [23] aim to solve this problem by attempting to fix a wide range of performance issues, using the same model instead of building an approach specialized to a specific category of issues. However, these approaches operate on isolated code snippets rather than full repository contexts that modern agents must handle.
II-C Benchmarks for Code Generation and Repair
@@
The evolution of coding benchmarks has progressed from simple algorithmic tasks to complex real-world scenarios. Early benchmarks like HumanEval [24] focused on isolated function generation, achieving high success rates (85-95%) but limited real-world applicability. The introduction of SWE-bench [9] marked a significant advancement by evaluating agents on real GitHub issues. Current state-of-the-art systems achieve varying performance: Claude models reach >50% on SWE-bench Verified, while more complex systems like AutoCodeRover achieve 18.83% on the full benchmark [25]. However, contamination concerns led to SWE-bench+ [26], where performance drops dramatically to 0.55-12%, revealing the importance of rigorous evaluation.
@@
Recent specialized benchmarks have emerged for specific domains. DevBench [27] evaluates comprehensive software development capabilities, while BigCodeBench [28] focuses on complex programming tasks requiring library usage. RefactorBench  [29] focuses on hand-crafted refactoring tasks in python repositories. However, none specifically target performance optimization tasks.
II-D Language Model Agents for Software Engineering
@@
Recent advances in coding agents have demonstrated impressive capabilities across software engineering tasks. SWE-agent [7] introduced specialized Agent-Computer Interfaces that significantly improved performance on repository-level tasks. OpenHands [8] provides a comprehensive platform achieving 29% success on SWE-bench Full and >60% on SWE-bench Verified with certain models. Several proprietary systems like Claude Code [3], Copilot Agent [4], Windsurf [5], Devin [6], etc. have been proposed and demonstrate impressive capabilities in fixing functional bugs. Recent work by Xia et al. [30] challenges the necessity of complex agent architectures, showing that simple three-phase approaches can achieve competitive performance (27.3%) at dramatically lower costs. This suggests that for specialized tasks like performance optimization, targeted approaches may be more effective than general-purpose agent frameworks. However, current agent evaluation focuses almost exclusively on functional bugs when it comes to bug-fixing. We believe that performance bugs require fundamentally different reasoning patterns, including understanding computational complexity, resource utilization, and performance trade-offs. Agents must also master performance-specific tooling and benchmarking methodologies to validate improvements, capabilities that have not been systematically evaluated.
II-E Performance Optimization and LLM Applications
@@
Limited work exists in applying LLMs to performance optimization. Traditional performance optimization relies heavily on profiling tools and expert knowledge. Frameworks like BenchmarkDotNet [31] for .NET provide comprehensive performance measurement capabilities, but require significant expertise to interpret results and guide optimization efforts. Some recent work has explored LLM-based approaches to performance issues. DeepDev-PERF [22] demonstrates a deep learning-based approach for improving software performance, while PerfLens [32] provides a data-driven performance bug detection and fix platform. However, these approaches operate on isolated methods rather than full repository contexts that modern agents must handle. Research has also developed retrieval-augmented prompt generation approaches for performance bug fixing. Recent work demonstrates that leveraging knowledge-bases of past performance fixes can improve LLM-generated solutions through targeted prompt engineering [33, 23]. However, this work operates in controlled settings with pre-identified performance issues rather than the complex repository navigation and issue identification, tasks that real-world agents must perform. Our work bridges these areas by creating an automated framework where agents can generate and validate performance improvements through self-designed benchmarks in realistic repository contexts, addressing the gap between isolated performance optimization and comprehensive software engineering agent evaluation for repo-level improvements.
III PerfBench Construction
@@
III-A Data Collection Process
We collected performance bug fixes through a systematic process. We first identified $\sim$1200 .NET repositories on GitHub with >10 stars, focusing on actively maintained projects with substantial codebases. Within the selected repositories, we crawl the commit history and find commits containing performance-related keywords: "performance", "slow", "optimization", "memory", "CPU", "latency", "throughput", etc. A complete list of all the keywords is provided in the appendix. We then identified commits that reference an issue and modify at least one .cs file i.e. code changes excluding documentation or configuration changes. For each candidate repo, we create a Docker environment by installing the appropriate version of .NET, available in the .csproj files and verify that both the buggy and fixed versions compile successfully using the dotnet build command. Finally, two experienced .NET developers independently reviewed each issue description and associated fix and issue to confirm that it represents a genuine performance bug and improvement. This process yielded 81 high-quality tasks spanning diverse set of repositories and performance issue types. Figure 1 shows the distribution of repos within our benchmark. We can see that unlike benchmarks like SWE-Bench [9], our benchmark isn’t concentrated within a small set of popular python repos. Instead we draw examples from 32 different repos from various domains.
Figure 1: Distribution of C# repositories in the benchmark.
@@ III-A Data Collection Process
Figure 2: Example task from PerfBench showing the metadata collected in our benchmark. The task given to the agent includes repository context and the original GitHub issue describing an OutOfMemoryException.
III-B Extracted Metadata
Each example within PerfBench contains a specific set of metadata collected to provide agents with the necessary context for mimicking a real-world development scenario. The metadata collected for each example includes: the repository name, the performance issue description, the commit hashes for both the buggy and fixed versions of the code as well as the ground truth patch. Each task is setup to run in an isolated docker container. Figure 2 illustrates a representative task from our benchmark, showing the metadata collected.
III-C Task Structure
@@
III-D Repository and Commit Statistics
To better understand the characteristics of the collected tasks, we analyze the collected GitHub issues, repositories in which the performance issues occurs as well as the gold patch commit. Table I summarizes key statistics across the 81 tasks in PerfBench.
TABLE I: Summary statistics of GitHub issues, repositories and gold patch fix in PerfBench.
@@ III-E Performance Bug Taxonomy
Our analysis reveals five major categories of performance bugs, with memory management issues being the most prevalent, followed by concurrency problems, algorithmic inefficiencies, I/O performance issues, and build tool performance problems. Table II presents the complete taxonomy with counts for each category.
The distribution shows that memory-related issues account for over 40% of all performance bugs in our benchmark, highlighting the critical importance of proper memory management and low allocations in .NET applications. Concurrency and algorithmic inefficiency issues each represent approximately 17% of the benchmark, while I/O and build / test performance issues comprise the remaining cases. In our later analysis, we also report the relative performances of agent and model configurations on each of these categories.
TABLE II: Taxonomy of performance bug types in PerfBench
@@
IV Experimental Setup
IV-A Evaluation Harness Design
The PerfBench evaluation harness automates the entire testing process. It extracts benchmark tests generated by the agent itself as well as the code changes containing performance improvements suggested by the agent. Benchmarks for .NET apps are typically written using the BenchmarkDotNet [31] framework. Our harness expects the agents to have written benchmark tests using the standard framework instead of rolling its own benchmarking code. This is a standard practice followed by performance engineers in .NET. We then execute the agent-written BenchmarkDotNet tests before and after the changes suggested by the agent and developer along with executing unit tests present within the repo. We report the success rate as well as several other metrics described below.
@@ IV-A Evaluation Harness Design
Figure 3: A sample BenchmarkDotNet output table showing execution time statistics, memory allocation and Garbage Collection (GC) metrics across different tests within the test suite. Our evaluation harness uses this to determine whether the code changes improve performance.
IV-B Metrics
For each task created above, the harness computes the following metrics:
@@ IV-B Metrics
  • Steps Taken (# steps): We also report the number of steps taken by the agent to solve the task. Steps taken translate to latency, so we want this to be low as well for high responsiveness.
@@
IV-C Agent Configurations
We evaluate the open-source Software Engineering agent OpenHands [8] against the benchmark in two different configurations:

### v3

@@ Abstract.
Performance bugs are inefficiencies in software that waste computational resources without causing functional failures, making them particularly challenging to detect and fix. While recent advances in Software Engineering agents have shown promise in automated bug fixing, existing benchmarks primarily focus on functional correctness and fail to evaluate agents’ abilities to identify and resolve non-functional issues like performance bugs. We introduce PerfBench, a benchmark comprising 81 real-world performance bug-fixing tasks from popular .NET repositories on GitHub. Unlike existing benchmarks that rely on pre-existing test suites, PerfBench features a novel evaluation harness that allows agents to generate their own performance benchmarks and validates fixes by comparing execution metrics collected for developer fix and agent fix. Each task in PerfBench is derived from actual developer fixes linked to performance-related issues, which are then verified by human experts, ensuring real-world relevance. Our evaluation reveals that current state-of-the-art coding agents struggle with performance optimization tasks, with baseline OpenHands agent achieving only a $\sim$3% success rate on our benchmark. We develop OpenHands-Perf-Agent, which incorporates performance-aware tooling and instructions and achieves a $\sim$20% success rate on the benchmark. We show that by ensuring the agent has proper instructions to benchmark its changes and tooling for benchmark output processing, we can improve the agent performance significantly, but room for improvement still remains. PerfBench provides a challenging test set for furthering the capabilities of agents in fixing perf issues.
1. Introduction
@@
Performance bugs represent a unique class of software defects that impact the application’s efficiency without causing any functional failures. Unlike traditional bugs that manifest as crashes or incorrect outputs, performance bugs silently waste computational resources, increase latency as well as costs (Attariyan et al., 2012; Dean et al., 2014a). These inefficiencies are particularly problematic in cloud applications where resource consumption directly translates to computational costs, and can also impact end-user experience due to problems such as increased latency.
@@
The rise of Software Engineering agents has revolutionized automated software engineering, with proprietary systems like Claude Code (Anthropic, 2024), Copilot Agent (GitHub, 2024), Windsurf (Windsurf, 2024), Devin (Cognition.ai, 2024), etc. as well as open-source agents such as SWE-agent (Yang et al., 2024), and OpenHands (Wang et al., 2024) demonstrating impressive capabilities in fixing functional bugs as well as other software engineering tasks such as test generation, code search, feature development, etc. However, current bug fixing benchmarks for evaluating these agents, such as SWE-bench (Jimenez et al., 2024), focus exclusively on functional correctness. This leaves a significant gap in understanding how well these agents can handle non-functional bugs such as performance or security bugs.
@@
2. Background and Related Work
@@
We discuss work related to performance bug detection and repair, benchmarks for code generation, and software engineering agents.
2.1. Performance Bugs in Software Systems
@@
Performance bugs represent a unique class of software defects that impact efficiency without causing functional failures. They tend to be harder to detect  (Attariyan et al., 2012; Dean et al., 2014a, b; Jovic et al., 2011) and fix  (Nistor et al., 2013; Song and Lu, 2014) than functional bugs. Due to being hard to detect and not causing outright failures, these bugs can often go undetected for long periods of time  (Dean et al., 2014b; Jovic et al., 2011). As a result, better tool support is needed to fix performance bugs. While performance profiling have been developed to identify these issues (Dean et al., 2014b; Attariyan et al., 2012; Han et al., 2012), these bugs require significant manual analysis to translate findings into fixes.
2.2. Automated Program Repair for Performance Issues
@@
Traditional automated program repair (APR) techniques focus primarily on functional correctness using test suites or logic assertions as specifications. GenProg (Weimer et al., 2009) uses genetic programming to generate fixes that pass test cases, while pattern-based approaches like PAR (Kim et al., 2013) leverage manually created fix templates. Similarly, CapGen (Wen et al., 2018) generates patches at finer granularity using context-aware prioritization, and VarFix (Wong et al., 2021) extends GenProg through variational execution (Austin and Flanagan, 2012). However, these traditional APR approaches face significant challenges when applied to performance bugs. Performance issues require different validation mechanisms than functional correctness, as they must demonstrate measurable improvements in efficiency metrics rather than passing test cases. Additionally, performance bugs often involve subtle inefficiencies that require domain-specific knowledge about algorithms, data structures, and system behavior that is difficult to encode in traditional search-based approaches. Not only that, performance bugs manifest in various forms including algorithmic inefficiencies, memory leaks, excessive allocations, and suboptimal API usage patterns. As a result, for a fix approach to be useful, it would need to be general enough to fix a wide-range of issues. However, many existing approaches target specific kinds of issues such as repeated computations (Della Toffola et al., 2015), software misconfigurations (Iqbal et al., 2021), loop inefficiencies (Nistor et al., 2013). Recent work in model training has explored performance-specific repair. Studies like DeepDev-PERF (Garg et al., 2022), RAPGen (Garg et al., 2025) aim to solve this problem by attempting to fix a wide range of performance issues, using the same model instead of building an approach specialized to a specific category of issues. However, these approaches operate on isolated code snippets rather than full repository contexts that modern agents must handle.
2.3. Benchmarks for Code Generation and Repair
@@
The evolution of coding benchmarks has progressed from simple algorithmic tasks to complex real-world scenarios. Early benchmarks like HumanEval (Chen et al., 2021) focused on isolated function generation, achieving high success rates (85-95%) but limited real-world applicability. The introduction of SWE-bench (Jimenez et al., 2024) marked a significant advancement by evaluating agents on real GitHub issues. Current state-of-the-art systems achieve varying performance: Claude models reach >50% on SWE-bench Verified, while more complex systems like AutoCodeRover achieve 18.83% on the full benchmark (Zhang et al., 2024). However, contamination concerns led to SWE-bench+ (Aleithan et al., 2024), where performance drops dramatically to 0.55-12%, revealing the importance of rigorous evaluation.
@@
Recent specialized benchmarks have emerged for specific domains. DevBench (Li et al., 2024) evaluates comprehensive software development capabilities, while BigCodeBench (Zhuo et al., 2025) focuses on complex programming tasks requiring library usage. RefactorBench  (Gautam et al., 2025) focuses on hand-crafted refactoring tasks in python repositories. However, none specifically target performance optimization tasks.
2.4. Language Model Agents for Software Engineering
@@
Recent advances in coding agents have demonstrated impressive capabilities across software engineering tasks. SWE-agent (Yang et al., 2024) introduced specialized Agent-Computer Interfaces that significantly improved performance on repository-level tasks. OpenHands (Wang et al., 2024) provides a comprehensive platform achieving 29% success on SWE-bench Full and >60% on SWE-bench Verified with certain models. Several proprietary systems like Claude Code (Anthropic, 2024), Copilot Agent (GitHub, 2024), Windsurf (Windsurf, 2024), Devin (Cognition.ai, 2024), etc. have been proposed and demonstrate impressive capabilities in fixing functional bugs. Recent work by Xia et al. (Xia et al., 2024) challenges the necessity of complex agent architectures, showing that simple three-phase approaches can achieve competitive performance (27.3%) at dramatically lower costs. This suggests that for specialized tasks like performance optimization, targeted approaches may be more effective than general-purpose agent frameworks. However, current agent evaluation focuses almost exclusively on functional bugs when it comes to bug-fixing. We believe that performance bugs require fundamentally different reasoning patterns, including understanding computational complexity, resource utilization, and performance trade-offs. Agents must also master performance-specific tooling and benchmarking methodologies to validate improvements, capabilities that have not been systematically evaluated.
2.5. Performance Optimization and LLM Applications
@@
Limited work exists in applying LLMs to performance optimization. Traditional performance optimization relies heavily on profiling tools and expert knowledge. Frameworks like BenchmarkDotNet (.NET Foundation, 2024) for .NET provide comprehensive performance measurement capabilities, but require significant expertise to interpret results and guide optimization efforts. Some recent work has explored LLM-based approaches to performance issues. DeepDev-PERF (Garg et al., 2022) demonstrates a deep learning-based approach for improving software performance, while PerfLens (Garg et al., 2021) provides a data-driven performance bug detection and fix platform. However, these approaches operate on isolated methods rather than full repository contexts that modern agents must handle. Research has also developed retrieval-augmented prompt generation approaches for performance bug fixing. Recent work demonstrates that leveraging knowledge-bases of past performance fixes can improve LLM-generated solutions through targeted prompt engineering (Liu et al., 2023; Garg et al., 2025). However, this work operates in controlled settings with pre-identified performance issues rather than the complex repository navigation and issue identification, tasks that real-world agents must perform. Our work bridges these areas by creating an automated framework where agents can generate and validate performance improvements through self-designed benchmarks in realistic repository contexts, addressing the gap between isolated performance optimization and comprehensive software engineering agent evaluation for repo-level improvements.
3. PerfBench Construction
@@
3.1. Data Collection Process
We collected performance bug fixes through a systematic process. We first identified $\sim$1200 .NET repositories on GitHub with >10 stars, focusing on actively maintained projects with substantial codebases. Within the selected repositories, we crawl the commit history and find commits containing performance-related keywords: "performance", "slow", "optimization", "memory", "CPU", "latency", "throughput", etc. A complete list of all the keywords is provided in Table 1. We then identified commits that reference an issue and modify at least one .cs file i.e. code changes excluding documentation or configuration changes. For each candidate repo, we create a Docker environment by installing the appropriate version of .NET, available in the .csproj files and verify that both the buggy and fixed versions compile successfully using the dotnet build command. Finally, two experienced .NET developers independently reviewed each issue description and associated fix and issue to confirm that it represents a genuine performance bug and improvement. This process yielded 81 high-quality tasks spanning diverse set of repositories and performance issue types. Figure 1 shows the distribution of repos within our benchmark. We can see that unlike benchmarks like SWE-Bench (Jimenez et al., 2024), our benchmark isn’t concentrated within a small set of popular python repos. Instead we draw examples from 32 different repos from various domains.
Table 1. Performance-Related Keywords Used for Initial Issue Filtering
Category Keywords
Core Performance perf, performance, optimize, optimization, faster, slower, speed up, slow down, latency, throughput, overhead, efficiency, scalable, scalability, bottleneck, lag, load time
Memory & CPU memory, alloc, allocation, dealloc, leak, heap, stack, gc, garbage collection, cpu, utilization, cache, cache miss, oom, out of memory, boxing
Execution & Runtime hot path, critical path
User-Perceived Issues hang, freeze, unresponsive, laggy, delay
Performance Signals inefficient, excessive, high cpu, high memory, not scalable, timing issue
Figure 1. Distribution of C# repositories in the benchmark.
@@ 3.1. Data Collection Process
Figure 2. Example task from PerfBench showing the metadata collected in our benchmark. The task given to the agent includes repository context and the original GitHub issue describing an OutOfMemoryException.
3.2. Extracted Metadata
Each example within PerfBench contains a specific set of metadata collected to provide agents with the necessary context for mimicking a real-world development scenario. The metadata collected for each example includes: the repo name, the performance issue description, the commit hashes for the buggy and fixed versions of the code as well as the ground truth patch. Each task is setup to run in an isolated docker container. Figure 2 illustrates a representative task from our benchmark, showing the metadata collected.
3.3. Task Structure
@@
3.4. Repository and Commit Statistics
To better understand the characteristics of the collected tasks, we analyze the collected GitHub issues, repositories in which the performance issues occurs as well as the gold patch commit. Table 2 summarizes key statistics across the 81 tasks in PerfBench.
Table 2. Summary statistics of GitHub issues, repositories and gold patch fix in PerfBench.
@@ 3.5. Performance Bug Taxonomy
Our analysis reveals five major categories of performance bugs, with memory management issues being the most prevalent, followed by concurrency problems, algorithmic inefficiencies, I/O performance issues, and build tool performance problems. Table 3 presents the complete taxonomy with counts for each category.
The distribution shows that memory-related issues account for >40% of all performance bugs in our benchmark, highlighting the critical importance of proper memory management and low allocations in .NET applications. Concurrency and algorithmic inefficiency issues each represent approximately 17% of the benchmark, while I/O and build / test performance issues comprise the remaining cases. In our later analysis, we report the relative performances of agent and model configurations on each category.
Table 3. Taxonomy of performance bug types in PerfBench
@@
4. Experimental Setup
4.1. Evaluation Harness Design
The PerfBench evaluation harness automates the entire testing process. It extracts benchmark tests generated by the agent itself as well as the code changes containing performance improvements suggested by the agent. Benchmarks for .NET apps are typically written using the BenchmarkDotNet (.NET Foundation, 2024) framework. Our harness expects the agents to have written benchmark tests using the standard framework instead of rolling its own benchmarking code. This is a standard practice followed by performance engineers in .NET. We then execute the agent-written BenchmarkDotNet tests before and after the changes suggested by the agent and developer along with executing unit tests present within the repo. We report the success rate as well as several other metrics described below.
@@ 4.1. Evaluation Harness Design
Figure 3. A sample BenchmarkDotNet output table showing execution time statistics, memory allocation and Garbage Collection (GC) metrics across different tests within the test suite. Our evaluation harness uses this to determine whether the code changes improve performance.
4.2. Metrics
For each task, the harness computes the following metrics:
@@ 4.2. Metrics
  • Steps Taken (# steps): We also report the number of steps taken by the agent to solve the task. Steps translate to latency, so we want this to be low for high responsiveness.
@@
4.3. Agent Configurations
We evaluate the open-source Software Engineering agent OpenHands (Wang et al., 2024) against the benchmark in two different configurations: